OpenStudy (anonymous):

Please help, thanks! Juan deposited $200 in a savings account earning 3% interest over 2 years. What was the total amount of interest earned at the end of 2 years? A. $12 $12?

OpenStudy (anonymous):

**Middle school

OpenStudy (anonymous):

answers are: A. $12 B. $24 C. $120 D. $240

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Assuming this is simple interest, then... I = P*r*t I = 200*0.03*2 I = 12 so yes you earn $12 in simple interest